angle a is 4 times as large as angle b and angle c is 20 degrees less than angle b...what does each angle equal?
Answer by rapaljer(4551)

Since angle a is given in terms of angle b and angle c is given in terms of angle b, everything is in terms of angle b, so let x = angle b.

Let x = angle b
4x = angle a
x-20 = angle c

The sum of the angles is 180 degrees.
x + 4x + x - 20 = 180
6x - 20 = 180
6x = 200 {Note: Answer doesn't come out even, but then that's life sometimes!)
x= 200/6 = 100/3 = 33 1/3 degrees -- angle b
4x = 400/3 = 133 1/3 degrees -- angle a
x-20 = 33 1/3 - 20 = 13 1/3 degress -- angle c
Check: 33 1/3 + 133 1/3 + 13 1/3 = 180 degrees
It checks!

R^2 at SCC

Check: